### v4

# Modelling and calibration of pair-rule protein patterns in Drosophila embryo: From Even-skipped and Fushi-tarazu to Wingless expression networks

Abstract
We modelled and calibrated the distributions of the seven-stripe patterns of Even-skipped (Eve) and Fushi-tarazu (Ftz) pair-rule proteins along the anteroposterior axis of the Drosphila embryo, established during early development. We have identified the putative repressive combinations for five Eve enhancers, and we have explored the relationship between Eve and Ftz for complementary patterns. The regulators of Eve and Ftz are stripe-specific DNA enhancers with embryo position-dependent activation rates and are regulated by the gap family of proteins. We achieved remarkable data matching of the Eve stripe pattern, and the calibrated model reproduces gap gene mutation experiments. Extended work inferring the Wingless (Wg) fourteen stripe pattern from Eve and Ftz enhancers have been proposed, clarifying the hierarchical structure of Drosphila’s genetic expression network during early development.
keywords:
Drosophila melanogaster early development , Genetic segmentation network, Stripe enhancers of Eve and Ftz , Segment-polarity protein regulators
1 Introduction
The emergence of periodic band structures along the anteroposterior axis of the Drosophila embryo during early development characterises Drosophila embryogenesis, one of the most thoroughly studied gene regulatory networks (Nüsslein-Volhard, (2006)). Broad domains of maternal proteins regulate the expression of a set of zygotic genes – the gap genes, along the anteroposterior axis of the embryo (Akam, (1997), Rivera-Pomar et al., (1996), Alves et al., (2006)). At a later stage of development, a new set of proteins form seven-stripe complementary patterns – pair-rule proteins, of which the most studied pair is the Even-skipped (Eve) and Fushi-tarazu (Ftz) (Frash et al., (1987)). This cascade of gene-expressed proteins (maternal $\to$ gap $\to$ pair-rule) terminates with the segment-polarity gene-expressed proteins that double the pattern to fourteen stripes, after which cellularisation begins. Cells in the region of each strip, at a later stage of development, will determine the segmented larvae body plan.
It is presumed that each Eve stripe, and possibly each Ftz stripe, is formed and controlled individually by a specific set of transcription factors that regulate individual enhancers (Nüsslein-Volhard, (2006)) – regions of the DNA that affect gene transcription and where transcription factors bind. Over the past decades, eve’s transcriptional control was studied via mutation experiments. The most recent depiction of the eve locus includes five separate enhancers: two dual stripe elements 3+7 and 4+6, and three single stripe elements 1, 2 and 5 (Lim et al., (2018)). The classical view of the stripe 2 enhancer includes two transcription repressors, Krüppel (Kr) and Giant (Gt), and two transcription activators, Bicoid (Bcd) and Hunchback (Hb), who work synergistically (Small et al., (1992), Simpson-Brose et al., (1994)). However, Vincent et al., (2018) proposed an alternative hypothesis that Hb represses the stripe 2 enhancer and that its repression can be counteracted by nearby sequences, possibly by Caudal (Cad), a protein whose mRNA has maternal origin. As for stripe 5, little information is provided, but it is proposed to be regulated by two transcription repressors, Kr and Gt (Fujioka et al., (1999)). Finally, the two dual stripe enhancers are believed to be regulated by the same transcription repressors, Hb and Knirps (Kni) (Small et al., (1996), Fujioka et al., (1999)). In the latter case, it is suggested that the enhancers are sensitive to repression by the number and affinity to repressor-binding sites, implying that the stripes 3+7 and 4+6 enhancers would respond autonomously to different amounts of Hb and Kni (Clyde et al., (2003), Crocker et al., (2016)). It is also suggested that although stripe 7 is not separable from stripe 3, its activation requires sequences outside of the minimal stripe 3 element. This led to the suggestion of an extra enhancer, a 2+7 dual element, which would be an extended version of the minimal eve stripe 2 enhancer (Vincent et al., (2018)). To date, practically nothing is known about the formation of stripe 1, with only a mention that it may be activated by Hb (Fujioka et al., (1999)).
Despite the many experiments for Eve, there is still no consensus on the minimum set of its regulators or any successful mathematical modelling, calibration, or spatial distribution of the full seven-stripe pattern of Eve. It is even less clear how the formation of the Ftz pattern occurs. The latest hypothesis (Schroeder et al., (2011)) suggests three enhancers initiating stripe pairs 1+5, 2+7, and 3+6, and the Zebra enhancer driving the expression of stripe 4, but, as far as we know, no transcription factors are known or reported in the literature.
In this paper, we develop a simple kinetic framework that accurately predicts and calibrates the formation of the Eve stripe pattern, providing a minimum set of gap regulators for each eve enhancer. We use the mass-action law to describe every biological process describing the expression of Eve along the anteroposterior axis of the Drosophila embryo. We anchor this work on the data provided by the FlyEx database (Kozlov et al., (2000), Myasnikova et al., (2001), Poustelnikova et al., (2004), Pisarev et al., (2008)). Based on the success of this modelling approach for Eve, we can predict the position of six of the seven stripes of Ftz, and we have extended our approach to predict the fourteen stripes of the segment-polarity protein Wingless (Wg). The nomenclature for gene and protein names follows the convection adapted from (Alberts et al.,, 2022, p. ix): the first letter of protein names is uppercase, and the acronymous follows the same rule but is typed in italics. For genes, the first letter is lowercase and typed in italics.
2 Methods
To describe all the genetic regulatory interactions that occur during the embryonic segmentation of Drosophila, we adopted the Alves et al., (2005) approach. We considered that the kinetic mechanisms describing the non-regulated production of a protein P is
$\begin{array}[]{l}\displaystyle\text{G}\mathop{\longrightarrow}^{k_{1}}\text{G}+\text{P}\\ \displaystyle\text{P}\mathop{\longrightarrow}^{d}\end{array}$ (1)
where G is the gene for P. We assume that DNA splicing has occurred, the transcription and translation from DNA to protein and the concentration of polymerases are represented by the rate constant $k_{1}$, and that $d$ is the intrinsic degradation rate of protein P. The generic kinetic mechanisms that describe gene repression is
$\begin{array}[]{l}\displaystyle\text{G}\mathop{\longrightarrow}^{k_{1}}\text{G}+\text{P}\\ \displaystyle\text{R}+\text{G}\mathop{\rightleftarrows}^{k_{2}}_{k_{-2}}\text{G}_{\text{R}}\\ \displaystyle\text{P}\mathop{\longrightarrow}^{d}\end{array}$ (2)
[…]
We tested models with different combinations of repressive and activating transcription factors, but those originating the observed experimental distribution of Eve are presented in Figure 1. The regulation of the Eve stripes as depicted in Figure 1 is obtained by several combinations of the basic mechanisms (1)-(2). The choice of wild-type eve locus has been adapted from Lim et al., (2018). We have added two more repressor locus sites for Tll and Hkb, defining the anterior and posterior regions of the Drosophila’s embryo (Weigel et al., (1990) and Casanova, (1990)).
The complete list of kinetic mechanisms specific to each enhancer locus for eve and ftz is presented in the Appendices. We applied the mass-action law to each kinetic mechanism and obtained the differential equations describing the expression of the corresponding Eve and Ftz stripes along the anteroposterior axis of the Drosophila’s embryo. The differential equation models to be calibrated were derived using the software package Kinetics (Dilão et al., (2010)) written for Mathematica programming language.
Figure 1: Schematic of the wild-type eve locus with two dual stripe elements 3+7 and 4+6 and three single stripe enhancers 1, 2 and 5. Adapted from Lim et al., (2018). Two repressor locus sites for Tailless (Tll) and Huckebein (Hkb) were added, defining the anterior and posterior regions of the Drsophila’s embryo where Eve expression is absent at steady state.
We have not considered diffusion mechanisms for proteins or mRNA in the present paper modelling approach. Reaction-diffusion-type models are used to describe the formation of maternal and gap gene families of proteins (Alves et al., (2006), Dilão et al., 2010b , Dilão, (2014)). Here, we have assumed that the stripes of the pair-rule proteins are formed due to the action of repressive transcription factors distributed along the anteroposterior axis of the Drosophila’s embryo, with a clear steady-state expression during the fourteenth mitotic cycle of the Drosphila zygote.
Figure 2: Distribution of gap protein (experimental – dots – and filtered) along the anteroposterior axis of the Drosophila’s embryo from cycle 14A4. Ordinate coordinates for each protein are arbitrary and do not correlate. FlyEx datasets: pe3 (Kr, Gt), rb9 (Hb, Kni) and th4 (Tll). Data for Hkb is from the SuperFly database: batch 220310, embryo 016. The embryo lengths were rescaled to $L=1$.
As the distributions of the concentrations of the different proteins along the Drosophila’s embryo are not measured in well-defined units, the model equations were adapted so that the variables of interest in the model incorporate a calibration parameter $\alpha_{P}$, specific to each protein P. For each protein P, we introduce a new variable
$=\alpha_{\text{p}}\text{P}$ P (3)
where $\alpha_{\text{p}}$ is the calibration parameter of a constituent in each reaction. Introducing the calibration parameter $\alpha_{P}$ is necessary because protein concentrations are measured in arbitrary units and differ for the plethora of experiments. Then, the differential equations in time were solved numerically, with each function taking the experimental spatial distribution of the modelled proteins as input, symbolised by $\overline{\text{P}}$. Model equations with the respective calibration parameters and remaining parameters are listed in the Appendices.
To determine the set of parameters that produce a calibrated steady-state solution that best explains the spatial distribution of Eve, we implemented a Monte Carlo algorithm where, at each iteration, every calibration parameter $\alpha_{i}$, reaction rate $k_{i}$ and $d$ were generated randomly in the spatial strip domains of the zygote. A fitness function evaluates the quality of the steady-state solutions of the model equations when compared with the data. We represent by $B_{exp}(x_{i})$, with $i=1,\ldots,n$, the distribution of the experimental data points along the embryo, where $x_{1}=0$ and $x_{n}=L=1$ are the anterior and posterior positions of the embryo, respectively. The resulting steady-state distribution of a given protein derived from the model is represented by $B(x_{i};\vec{p})$, where $\vec{p}=(p_{1},...,p_{m})$ is the set of $m$ simulation parameters. The fitness function is defined as the average of the squared residuals
$\chi^{2}(\vec{p})=\frac{1}{n}\sum_{i=1}^{n}\left(B(x_{i};\vec{p})-B_{exp}(x_{i})\right)^{2},$
and the quality of the fit is measured by $\sqrt{\chi^{2}(\vec{p})}$. The relative error of the fit of stripe $i$ fit is estimated by $\varepsilon_{i}=\sqrt{\chi^{2}_{i}/B^{2}(x_{i};\vec{p})}$, where $x_{i}$ is the position where the stripe $i$ is maximum (Materials and methods in Dilão et al., 2010c ).
After $M$ iterations of the Monte Carlo algorithm, the set of parameters providing the lower $\chi^{2}$ value was chosen and used to deliver the best fit to the experimental data of each model.
All model results correspond to the best fit out of half a million trials with Monte Carlo optimisation ($M=0.5\times 10^{6}$).
Data was taken from the FlyEx database (Kozlov et al., (2000), Myasnikova et al., (2001), Poustelnikova et al., (2004), Pisarev et al., (2008)), except for Hkb, which was taken from the SuperFly database (Cicin-Sain et al., (2014)). Following common data processing practices (Surkova et al., (2008)), we used a Laplacian filter to remove background noise and levelled all distributions to eliminate residual fluorescent concentrations (Alves-Pires et al., (1998)). The size of the embryos was rescaled to L=1. The final gap protein profiles along the normalised anteroposterior position in the embryo of Drosophila and the corresponding data are shown in Figure 2.
To obtain the transcriptional regulators for Eve’s stripes, we analysed the distribution of gap proteins along Drosphila’s embryo (Figure 3). The observed distributions suggested that the gap transcription factors act as repressors of eve and work together combinatorially. Therefore, we adopted the concept of individual eve enhancers, which only allow Eve’s production when none of its repressive regulators binds to it, as expressed by mechanism (2). Assuming a kinetic system where the protein Eve is freely produced – mechanism (1) with a gene $\text{G}\to\text{G}_{e}$ – and has a subset of repressive gap protein regulators R, each repressor contributes to Eve’s production and its anteroposterior differentiated distribution. Higher production of Eve will occur in regions with lower sums of repressors’ concentration, resulting in a localised stripe formation. We tested all the possible combinations of two, three, and four gap proteins, assigning each protein a calibration parameter factor to mimic the calibration parameter $\alpha_{P}$, and obtained minimal sets of repressive transcriptional regulators.
Figure 3: Distribution of Eve’s stripes and gap proteins along the Drosphila’s embryo. The gap data are from stage 14A4, and Eve’s data is from stage 14A8. Ordinate coordinates for each protein are arbitrary and do not correlate.
[…]
3 Results and Discussion
In the Drosophila wild-type embryo, the pair-rule stripes are limited to a specific region of the Drosophila embryo, and Eve is not present at the anterior and posterior poles. We considered that the Eve stripe region is well-determined by the proteins (Tll) and (Hkb), only produced at the embryo ends. These proteins restrict the expression of Eve to the desired region by repressing its transcription at the poles. Therefore, the identified terminal regulators integrated every model developed to describe the formation of Eve’s stripes. Moreover, we considered only transcription factors of the gap family of proteins as there is evidence that maternal transcription factors at cell cycle 14 have no transcript activity (Winata et al., (2018)).
3.1 Eve’s stripes definition via a gap repressive system
$\mathbf{k_{eve-s1}}$ $\mathbf{k_{eve-s2}}$ $\mathbf{k_{eve-s3}}$ $\mathbf{k_{eve-s4}}$ $\mathbf{k_{eve-s5}}$ $\mathbf{k_{eve-s6}}$ $\mathbf{k_{eve-s7}}$
0.8584 0.8584 0.9500 3.1100 0.9150 0.8584 1.2900
Table 1: Eve production rate values for stripe 1-7 that best fit the experimental data for Eve at time class 8, cycle 14 of Drosophila embryogenesis shown in Figure 4.
$\mathbf{k_{kni}}$ $\mathbf{k_{-kni}}$ $\mathbf{\alpha_{kni}}$ $\mathbf{k_{hb}}$ $\mathbf{k_{-hb}}$ $\mathbf{\alpha_{hb}}$ d
1 2.8995 0.0002 0.5667 - - - 0.7906
2 5.8021 0.0304 0.9269 - - - 0.7906
3+7 4.0241 0.1620 0.1153 1.4049 0.0078 0.1523 0.7906
4+6 7.4764 0.0647 0.6720 7.1382 0.0424 0.0200 0.7906
5 - - - 3.6212 0.0657 0.0103 0.7906
$\mathbf{k_{kr}}$ $\mathbf{k_{-kr}}$ $\mathbf{\alpha_{kr}}$ $\mathbf{k_{gt}}$ $\mathbf{k_{-gt}}$ $\mathbf{\alpha_{gt}}$
1 3.3110 0.0001 0.0065 0.1294 0.0098 0.5990
2 0.7502 0.0048 0.0419 0.1091 0.0705 0.0819
3+7 - - - - - -
4+6 - - - 1.5716 0.0428 0.7840
5 7.4532 0.0356 0.5312 6.5580 0.0374 0.0894
$\mathbf{\sqrt{\chi^{2}_{1}}}$ $\mathbf{\sqrt{\chi^{2}_{2}}}$ $\mathbf{\sqrt{\chi^{2}_{3}}}$ $\mathbf{\sqrt{\chi^{2}_{4}}}$ $\mathbf{\sqrt{\chi^{2}_{5}}}$ $\mathbf{\sqrt{\chi^{2}_{6}}}$ $\mathbf{\sqrt{\chi^{2}_{7}}}$
4.1 3.2 13.2 36.1 4.3 17.8 19.6
$\mathbf{\varepsilon_{1}}$ $\mathbf{\varepsilon_{2}}$ $\mathbf{\varepsilon_{3}}$ $\mathbf{\varepsilon_{4}}$ $\mathbf{\varepsilon_{5}}$ $\mathbf{\varepsilon_{6}}$ $\mathbf{\varepsilon_{7}}$
0.07 0.05 0.11 0.33 0.07 0.27 0.33
Table 2: Parameter values that best fit the experimental data of Eve’s stripes at time class 8 of mitotic cycle 14 of Drosophila embryogenesis shown in Figure 4. Terminal gap repression parameters were fixed with the 3+7 simulation: $k_{tll}=2.2956$, $k_{-tll}=0.0747$, $\alpha_{tll}=0.7925$, $k_{hkb}=2.4229$, $k_{-hkb}=10^{-5}$, $\alpha_{hkb}=0.7988$. Model results after 500,000 trials with Monte Carlo, with calibration parameters $\alpha$ determined within the interval [0,1]. The quantity $\sqrt{\chi^{2}(\vec{p})}=\sqrt{\chi^{2}_{i}}$ measures the quality of the fit of stripe number $i$, and $\varepsilon_{i}=\sqrt{\chi^{2}_{i}/B^{2}(x_{i};\vec{p})}$ is an estimate of the relative error of each stripe fit.
In the first attempt of the model calibration, we considered that the production rate of Eve was the same across all cells. However, we found that adjusting Eve’s production spatially was necessary to improve the accuracy of the fits. For this reason, we considered the production rate as a function of space and adjusted it stripe by stripe to best match the data (Table 1). This approach is supported by recent studies demonstrating mRNA translation’s profound heterogeneity, including protein synthesis rates (Sonneveld et al., (2020)).
3.1.1 Single stripe 1 enhancer of Eve
To explain the formation of stripe 1 of the Eve protein, we propose a novel enhancer model regulated by the repressive proteins Tll, Hkb, Kr, Kni and Gt, with kinetic mechanisms
$\begin{array}[]{ll}\displaystyle\text{G}_{e}\mathop{\longrightarrow}^{k_{1}}\text{G}_{e}+\text{Eve},&\displaystyle\text{Eve}\mathop{\longrightarrow}^{d},\\ \displaystyle\text{Kr}+\text{G}_{e}\mathop{\rightleftarrows}^{k_{3}}_{k_{-3}}\text{G}_{R1},&\displaystyle\text{Kni}+\text{G}_{e}\mathop{\rightleftarrows}^{k_{4}}_{k_{-4}}\text{G}_{R2},\\ \displaystyle\text{Gt}+\text{G}_{e}\mathop{\rightleftarrows}^{k_{5}}_{k_{-5}}\text{G}_{R3},&\displaystyle\text{Tll}+\text{G}_{e}\mathop{\rightleftarrows}^{k_{6}}_{k_{-6}}\text{G}_{R4},\\ \displaystyle\text{Hkb}+\text{G}_{e}\mathop{\rightleftarrows}^{k_{7}}_{k_{-7}}\text{G}_{R5},&\end{array}$ (4)
where $\text{G}_{e}$ is the gene of Eve. The corresponding evolution equations derived from the mass action law are shown in A. The parameters of the model that best fit the data for stripe 1 are shown in Tables 1 and 2. The parameters in the kinetic mechanisms (4) in Tables 1 and 2 translate to: (Table 1) $k_{1}=k_{eve-s1}$, (Table 2) $k_{3}=k_{kr}$, $k_{-3}=k_{-kr}$, $k_{4}=k_{kni}$, $k_{-4}=k_{-kni}$, $k_{5}=k_{gt}$, $k_{-5}=k_{-gt}$, $k_{6}=k_{tll}$, $k_{-6}=k_{-tll}$, $k_{7}=k_{hkb}$ and $k_{-7}=k_{-hkb}$.
In Figure 4, we show the position and shape of stripe 1 along the anteroposterior axis of the Drosophila embryo, showing a remarkable agreement with the experimental data, reflected in a relative error of $\varepsilon_{1}=7\%$ (Table 2).
3.1.2 Single stripe 2 enhancer of Eve
When testing the joint formation of stripes 2 and 5 of Eve with only Gt and Kr, we realised that the only option was to separate their regulation with two single stripe enhancers. Thus, we propose Kr, Kni, and Gt to be the minimal gap set required to explain stripe 2, just as for stripe 1, with the same kinetic mechanisms (4), but different parameters. In this case, (Table 1) $k_{1}=k_{eve-s2}$ and the other parameters are defined as in the stripe 1 case.
With this assumption, we obtained a very precise prediction of stripe 2 – Figure 4, with a relative error lower than $\varepsilon_{2}=5\%$ (Table 2). Whereas for stripe 1 Gt and Kni had more or less the same level of repression, now we find a greater affinity to both Gt and Kr compared to Kni (Table 2).
3.1.3 Dual stripe enhancer 3+7 of Eve
We propose a dual stripe enhancer model for stripes 3 and 7 of Eve with the repressive regulators Tll, Hkb, Kni and Hb:
$\begin{array}[]{ll}\displaystyle\text{G}_{e}\mathop{\longrightarrow}^{k_{1}}\text{G}_{e}+\text{Eve},&\displaystyle\text{Eve}\mathop{\longrightarrow}^{d},\\ \displaystyle\text{Kni}+\text{G}_{e}\mathop{\rightleftarrows}^{k_{3}}_{k_{-3}}\text{G}_{R1},&\displaystyle\text{Hb}+\text{G}_{e}\mathop{\rightleftarrows}^{k_{4}}_{k_{-4}}\text{G}_{R2},\\ \displaystyle\text{Tll}+\text{G}_{e}\mathop{\rightleftarrows}^{k_{5}}_{k_{-5}}\text{G}_{R3},&\displaystyle\text{Hkb}+\text{G}_{e}\mathop{\rightleftarrows}^{k_{6}}_{k_{-6}}\text{G}_{R4}.\end{array}$ (5)
The equations associated with these kinetic mechanisms are shown in A.
The parameters best fitting the data for stripes 3 and 7 are shown in Tables 1 and 2. The parameters in the kinetic mechanisms (5) correspond to the following parameters in Tables 1 and 2: (Table 1) $k_{1}=k_{eve-s3}$, or $k_{1}=k_{eve-s7}$, or $k_{1}=0.8584$ away from the stripes 3 and 7 regions; (Table 2) $k_{3}=k_{kni}$, $k_{-3}=k_{-kni}$, $k_{4}=k_{hb}$, $k_{-4}=k_{-hb}$, $k_{5}=k_{tll}$, $k_{-5}=k_{-tll}$, $k_{6}=k_{hkb}$ and $k_{-6}=k_{-hkb}$.
The steady-state solution predicted – Figure 4 – shows the location, width, and amplitude of stripe 3 practically match the data, reflected in a low relative error of $\varepsilon_{3}=11\%$. Concerning stripe 7, we see that the anterior border forms correctly, but its location and width are not precise, with a high relative error of $\varepsilon_{7}=33\%$. This may be due to an excessive Hkb repression in the posterior embryo pole. The Hkb profile overlaps with stripe 7, suggesting that our model will always underestimate this stripe with the initial conditions data we had access to.
We showed that the proper functioning of this enhancer requires Kni to have a slightly higher repression calibration parameter, and at a faster rate, than Hb (Table 2).
Figure 4: Predictions for eve’s stripe-specific enhancers 1, 2, 5, 3+7 and 4+6 and full Eve pattern prediction obtained by combining the results of each enhancer model. The figure shows the best fits of each model to the experimental data of Eve stripes (dotted lines) at time class 8 of mitotic cycle 14 (14A8) of Drosophila embryogenesis. We obtained the parameter values of ki and $\alpha_{i}$ by a Monte Carlo optimisation after 500,000 trials, summarised in Table 2. The adjusted Eve production rates k1 are shown in Table 1.
3.1.4 Dual stripe enhancer 4+6 of Eve
For stripes 4 and 6 of the Eve pattern, we propose a model with the repressive regulators Tll, Hkb, Kni and Gt and Hb:
$\begin{array}[]{ll}\displaystyle\text{G}_{e}\mathop{\longrightarrow}^{k_{1}}\text{G}_{e}+\text{Eve},&\displaystyle\text{Eve}\mathop{\longrightarrow}^{d},\\ \text{Kni}+\text{G}_{e}\mathop{\rightleftarrows}^{k_{3}}_{k_{-3}}\text{G}_{R1},&\displaystyle\text{Gt}+\text{G}_{e}\mathop{\rightleftarrows}^{k_{4}}_{k_{-4}}\text{G}_{R2},\\ \displaystyle\text{Hb}+\text{G}_{e}\mathop{\rightleftarrows}^{k_{5}}_{k_{-5}}\text{G}_{R3},&\displaystyle\text{Tll}+\text{G}_{e}\mathop{\rightleftarrows}^{k_{6}}_{k_{-6}}\text{G}_{R4},\\ \displaystyle\text{Hkb}+\text{G}_{e}\mathop{\rightleftarrows}^{k_{7}}_{k_{-7}}\text{G}_{R5}.&\end{array}$ (6)
The equations associated with these kinetic mechanisms are shown in A. The parameters best fitting the data for stripes 4 and 6 are shown in Tables 1 and 2. The parameters in the kinetic mechanisms (6) correspond to the following parameters in Tables 1 and 2: (Table 1) $k_{1}=k_{eve-s4}$, or $k_{1}=k_{eve-s6}$, or $k_{1}=0.8584$ away from the stripes 4 and 6 regions; (Table 2) $k_{3}=k_{kni}$, $k_{-3}=k_{-kni}$, $k_{4}=k_{gt}$, $k_{-4}=k_{-gt}$, $k_{5}=k_{hb}$, $k_{-5}=k_{-hb}$, $k_{6}=k_{tll}$, $k_{-6}=k_{-tll}$, $k_{7}=k_{hkb}$ and $k_{-7}=k_{-hkb}$.
Regarding stripes 4 and 6, we achieved an overall correct prediction of both stripe’s localisation (Figure 4). However, localisation deviations of both stripes significantly increase their relative errors to around $\varepsilon_{4}=33\%$ and $\varepsilon_{6}=27\%$.
In this case, we found it necessary for Hb to have a much higher repressive contribution than Kni and Gt, contrary to the 3+7 case. This sensitivity difference was proposed in the literature (Clyde et al., (2003), Crocker et al., (2016)) and is now confirmed with our simple enhancer models. Furthermore, our results prove to be consistent with the computationally predicted number of binding sites for each enhancer: Clyde et al., (2003) identify 12 Hb and 16 Kni binding sites in the 3+7 enhancer, and 11 Hb and 4 Kni binding sites in the 4+6 enhancer. The proportions between these numbers are very similar to those between the calibration $\alpha^{\prime}s$ we computed: $\alpha_{Kni}=0.1152$ and $\alpha_{Hb}=0.1523$ for the 3+7 enhancer, and $\alpha_{Kni}=0.6720$ and $\alpha_{Hb}=0.0200$ for the 4+6 enhancer. These differences may be associated with the quality of different data sets.
3.1.5 Single stripe 5 enhancer of Eve
We found that Tll, Hkb, Kr, and Gt are sufficient to generate stripe 5 of Eve. However, when we added Hb as a repressive regulator, we improved the fit to the data, lowering the relative error from around $\varepsilon_{5}\simeq 13\%$ to less than $\varepsilon_{5}=8\%$ (Figure 4, Table 2). This latter enhancer has Hb as the main repressive regulator, followed by Gt, and both have much higher calibration parameter factors than Kr. Therefore, to regulate stripe 5, we have chosen the kinetic mechanism
$\begin{array}[]{ll}\displaystyle\text{G}_{e}\mathop{\longrightarrow}^{k_{1}}\text{G}_{e}+\text{Eve},&\displaystyle\text{Eve}\mathop{\longrightarrow}^{d},\\ \displaystyle\text{Kr}+\text{G}_{e}\mathop{\rightleftarrows}^{k_{3}}_{k_{-3}}\text{G}_{R1},&\displaystyle\text{Gt}+\text{G}_{e}\mathop{\rightleftarrows}^{k_{4}}_{k_{-4}}\text{G}_{R3},\\ \displaystyle\text{Hb}+\text{G}_{e}\mathop{\rightleftarrows}^{k_{5}}_{k_{-5}}\text{G}_{R4},&\displaystyle\text{Tll}+\text{G}_{e}\mathop{\rightleftarrows}^{k_{6}}_{k_{-6}}\text{G}_{R5},\\ \displaystyle\text{Hkb}+\text{G}_{e}\mathop{\rightleftarrows}^{k_{7}}_{k_{-7}}\text{G}_{R6}.&\end{array}$ (7)
The equations associated with these kinetic mechanisms are shown in A. The parameters that best fit the data for stripe 5 are shown in Tables 1 and 2. The parameters in the kinetic mechanisms (7) correspond to the following parameters in Tables 1 and 2: (Table 1) $k_{1}=k_{eve-s5}$, (Table 2) $k_{3}=k_{kr}$, $k_{-3}=k_{-kr}$, $k_{4}=k_{gt}$, $k_{-4}=k_{-gt}$, $k_{5}=k_{hb}$, $k_{-5}=k_{-hb}$, $k_{6}=k_{tll}$, $k_{-6}=k_{-tll}$, $k_{7}=k_{hkb}$ and $k_{-7}=k_{-hkb}$.
Overall, we accurately predicted all the stripes in the Eve pattern with a purely repressive gap regulating system. We found no justification or need for any gap transcription activator role. Moreover, our model does not require any maternal input. We propose that any effect on the Eve pattern observed in maternal gene mutation experiments comes from collateral damage to gap regulation and consequent disturbance in the formation of the gap domains.
3.2 Simulating gap mutations of Eve
An essential step in validating our model was the introduction of gap mutations. We removed one regulator at a time in each model and observed the effects on the Eve pattern of time class 8 of mitotic cycle 14 (Figure 5).
For a kr mutation introduced in stripe 2 (Figure 5, S2, kr-) and stripe 5 enhancers (Figure 5, S5, kr-), we observe the fusion of stripes 2 and 3 and the fusion of stripes 4 and 5. Such a result coincides with the experimental results of Surkova et al., (2013) and with the anterior expansion of stripe 5 reported in Fujioka et al., (1999). However, we do not observe the weakening of stripe 4 and the posterior expansion of stripe 6, reported in that same article, or even a more pronounced expression of stripe 2. We point out that we are not considering the collateral effects of each mutation in the expression of other gap genes. Thus, we disregard any indirect alteration of Eve’s pattern. For instance, a kr mutation can influence all other gap genes gt, kni and hb involved in forming stripes 4 and 6 (Alves et al., (2006)).
Figure 5: Effects on Eve’s pattern of the introduction of gap gene mutations $kr^{-}$, $gt^{-}$, $kni^{-}$ and $hb^{-}$ in each enhancer model, cycle 14A8 of Drosophila embryogenesis. A $kr^{-}$ embryo shows two fused domains, 2-3 and 4-5, coincident with the experimental results of Surkova et al., (2008) and Fujioka et al., (1999). A $gt^{-}$ embryo presents a merging of stripes 1 and 2 as of stripes 5 and 6, in agreement with Frash et al., (1987), Small et al., (1991) and Stanojevic et al., (1991). A $kni^{-}$ embryo presents a merging of stripes in the domain 3 to 7, just as observed in Surkova et al., (2008). A $hb^{-}$ embryo shows an anterior fused domain up to stripe 3 or 4, a posterior expansion of stripe 7, and a possible overlap with stripe 6. This is in line with the mutation experiments studied in Frash et al., (1987), Small et al., (1991) and Fujioka et al., (1999).
Next, we analysed the introduction of a gt mutation (Figure 5, S2 and S5, gt-), in which stripes 1 and 2 merge as well as stripes 5 and 6. These results agree with the experiments carried out by Frash et al., (1987) and also with Small et al., (1991) and Stanojevic et al., (1991) that report the expansion of the anterior border of stripe 2. Although an intensity reduction for stripe 7 is shown in Frash et al., (1987), we do not verify that. On the contrary, we observe the appearance of a seventh stripe that did not exist before on this model.
As for the kni mutation, it was introduced in both enhancers 3+7 and 4+6 (Figure 5, S3+7, kni- and S4+6, kni-). They show a fusion of stripes 3 to 7 and a fusion of stripes 4 to 6, the behaviour expected according to Surkova et al., (2013) experiments.
Finally, we studied the introduction of an hb mutation in the 3+7 and 4+6 enhancers (Figure 5, S3+7, hb- and S4+6, hb-). For the 3+7 case, we observe a broad fusion domain starting at the anterior end of the embryo and encompassing stripes 1, 2 and 3 and a posterior expansion of stripe 7. As for the 4+6 enhancer, we see a fused expression domain in the anterior region up to stripe 4 and an expansion of the posterior border of stripe 6, overlapping with stripe 7. Both cases are in line with the mutation experiences made by Frash et al., (1987), Small et al., (1996) and Fujioka et al., (1999) as they report a total or partial deletion of stripes 1 to 4, with a broad anterior band of expression instead, the expansion of stripes 3 and 7, and the expansion of stripes 4 and 6.
Both tll and hkb mutations were tested, but they did not cause significant changes to Eve’s pattern aside from a lack of definition in the embryo poles. Contrarily to what was expected (Small et al., (1996), Frash et al., (1987), Fujioka et al., (1999), Janssens et al., (2013), Ilsley et al., (2013)), a tll mutation did not abolish stripe 7. Once more, we believe that the abolition of this stripe is somehow related to the direct interference with the regulation of other gap genes, such as kni and gt, that is caused by a tll mutation. Furthermore, we hypothesise that the experimentally recorded effects of a mutation in maternal genes nanos (nos) and torso (tor) are once again the effect of the mis-expression of gap genes. For example, tor regulates the terminal gap genes, so a tor mutation is expected to alter Eve in a similar way to a mutation in tll or hkb.
3.3 Ftz’s stripes can be partially defined via a gap repressive system
Figure 6: Schematic of the hypothesised regulation of the wild-type ftz locus with three dual stripe elements 1+5, 2+7 and 3+6 stripe enhancers. Enhancers for stripe 4 are unknown.
Following a reasoning similar to the formation of Eve stripes, we tested the viability of stripe-specific enhancers driving Ftz stripes. The gap protein combinations we identified as possibly regulating Ftz correspond to those found to regulate eve’s transcription: our thermodynamics approach is compatible with two dual stripe elements 2+7 and 3+6 regulated by Hb and Kni and another dual stripe enhancer 1+5 regulated by Kr and Gt. We also assume that Tll and Hkb repress the production of Ftz in the anterior and posterior regions of the Drosophyla’s embryo (Figure 6).
Figure 7: Predictions for ftz’s stripe-specific enhancers 1+5, 2+7, 3+6 and full Ftz pattern prediction obtained by combining the results of each enhancer model. The figure shows the best fits of each model to the experimental data of Ftz stripes (dotted lines) at time class 8 of mitotic cycle 14 (14A8) of Drosophila’s embryogenesis. We obtained the parameter values of ki and $\alpha_{i}$ by a Monte Carlo optimisation after 500,000 trials, summarised in Table 4. The adjusted Ftz production rates k1 are listed in Table 3.
3.3.1 Dual stripe enhancer 1+5 of Ftz
We initially thought that Gt and Kr were enough to regulate the 1+5 enhancer, but we discovered that adding a third regulator, Kni, is essential. In this case, the kinetic regulatory mechanism is similar to the one for stripes 1 and 2 of Eve – kinetic mechanism (4) – with $\text{G}_{e}\to\text{G}_{f}$, where $\text{G}_{f}$ is the gene of Ftz. The corresponding evolution equations derived from the mass action law are shown in A. The parameters that best fit the data for stripes 1 and 5 are shown in Tables 3 and 4: (Table 3) $k_{1}=k_{ftz-s1}$, or $k_{1}=k_{ftz-s5}$, or $k_{1}=0.5823$ away from the stripes 1 and 5 regions; (Table 4) $k_{3}=k_{kr}$, $k_{-3}=k_{-kr}$, $k_{4}=k_{kni}$, $k_{-4}=k_{-kni}$, $k_{5}=k_{gt}$, $k_{-5}=k_{-gt}$, $k_{6}=k_{tll}$ and $k_{-6}=k_{-tll}$, $k_{7}=k_{hkb}$ and $k_{-7}=k_{-hkb}$.
In Figure 7, we show the position of stripes 1 and 5 along the anteroposterior axis of the Drosophila’s embryo, with a remarkable agreement with the experimental data, reflected in relative errors of $\varepsilon_{1}=20\%$ and $\varepsilon_{5}=18\%$, respectively (Table 2).
We found this enhancer to have a higher affinity to Kr and Kni than to Gt (Table 4).
$\mathbf{k_{ftz-s1}}$ $\mathbf{k_{ftz-s2}}$ $\mathbf{k_{ftz-s3}}$ $\mathbf{k_{ftz-s4}}$ $\mathbf{k_{ftz-s5}}$ $\mathbf{k_{ftz-s6}}$ $\mathbf{k_{ftz-s7}}$
1.3100 0.6800 1.8910 - 1.500 0.5630 3.3400
Table 3: Ftz production rate values that best fit the experimental data at time class 8 of mitotic cycle 14 of Drosophila embryogenesis shown in Figure 7.
$\mathbf{k_{kni}}$ $\mathbf{k_{-kni}}$ $\mathbf{\alpha_{kni}}$ $\mathbf{k_{hb}}$ $\mathbf{k_{-hb}}$ $\mathbf{\alpha_{hb}}$ d
1+5 5.6887 0.0205 0.9909 - - - 0.7104
2+7 7.3807 0.0034 0.0434 4.5349 0.0375 0.5605 0.7104
3+6 3.1028 0.0877 0.3160 7.7090 0.0265 0.0440 0.7104
$\mathbf{k_{kr}}$ $\mathbf{k_{-kr}}$ $\mathbf{\alpha_{kr}}$ $\mathbf{k_{gt}}$ $\mathbf{k_{-gt}}$ $\mathbf{\alpha_{gt}}$
1+5 3.6834 0.0748 0.2489 0.3609 0.0001 0.3461
2+7 - - - - - -
3+6 - - - - - -
$\mathbf{\sqrt{\chi^{2}_{1}}}$ $\mathbf{\sqrt{\chi^{2}_{2}}}$ $\mathbf{\sqrt{\chi^{2}_{3}}}$ $\mathbf{\sqrt{\chi^{2}_{4}}}$ $\mathbf{\sqrt{\chi^{2}_{5}}}$ $\mathbf{\sqrt{\chi^{2}_{6}}}$ $\mathbf{\sqrt{\chi^{2}_{7}}}$
14.2 32.4 10.3 - 14.1 15.9 18.0
$\mathbf{\varepsilon_{1}}$ $\mathbf{\varepsilon_{2}}$ $\mathbf{\varepsilon_{3}}$ $\mathbf{\varepsilon_{4}}$ $\mathbf{\varepsilon_{5}}$ $\mathbf{\varepsilon_{6}}$ $\mathbf{\varepsilon_{7}}$
0.20 0.36 0.09 - 0.18 0.19 0.18
Table 4: Parameter values that best fit the experimental data of Ftz’s stripes at time class 8 of mitotic cycle 14 of Drosophila’s embryogenesis shown in Figure 4. Terminal gap repression parameters fixed with the 2+7 simulation for Tll and with the 1+5 simulation for Hkb: $k_{tll}=5.2564$, $k_{-tll}=0.0937$, $\alpha_{tll}=0.4560$, $k_{hkb}=8.3229$, $k_{-hkb}=0.0500$, $\alpha_{hkb}=0.2900$. Model results after 500,000 trials with Monte Carlo, with calibration parameters $\alpha$ determined within the interval [0,1]. The quantity $\sqrt{\chi^{2}(\vec{p})}=\sqrt{\chi^{2}_{i}}$ measures the quality of the fit of stripe number $i$, and $\varepsilon_{i}=\sqrt{\chi^{2}_{i}/B^{2}(x_{i};\vec{p})}$ is an estimate of the relative error of each stripe fit.
3.3.2 Dual stripe enhancer 2+7 of Ftz
The position of stripes 2 and 7 are predicted with an enhancer model with repressive regulators Hb, Kni, Tll and Hkb. In this case, the kinetic regulatory mechanism is similar to the one for dual stripes 3+7 of Eve, (5), with $\text{G}_{e}\to\text{G}_{f}$. The corresponding evolution equations derived from the mass action law are shown in A. The parameters that best fit the data for stripes 2 and 7 are shown in Tables 3 and 4: (Table 3) $k_{1}=k_{ftz-s2}$, or $k_{1}=k_{ftz-s7}$, or $k_{1}=0.5823$ away from the stripes 2 and 7 regions; (Table 4) $k_{3}=k_{kni}$, $k_{-3}=k_{-kni}$, $k_{4}=k_{hb}$, $k_{-4}=k_{-hb}$, $k_{5}=k_{tll}$ and $k_{-5}=k_{-tll}$, $k_{6}=k_{hkb}$ and $k_{-6}=k_{-hkb}$.
An enhancer model with repressive regulators Hb and Kni predicted stripe 7 with a reasonable agreement with the data and a low relative error of about $\varepsilon_{7}=18\%$. Its location was precise, but its width was underestimated. As for stripe 2, it has a well-described anterior border but an extended posterior border. Such can be an indication that the Hb repression calibration parameter determined ($\alpha_{Hb}=0.5605$) is not enough. Together with the width mismatch, this leads to a high relative error of $\varepsilon_{2}=36\%$.
The proper functioning of this enhancer requires a higher repression calibration parameter from Kni versus Hb (Table 4).
3.3.3 Dual stripe enhancer 3+6 of Ftz
Finally, we studied the control mechanisms of the dual stripe enhancer 3+6, which we propose to be regulated by the same proteins as the 2+7 enhancer of Ftz and the 3+7 enhancer of Eve: Hb, Kni, Tll and Hkb. The predicted stripes match the experimental data, reflected in the low relative errors for stripes 3 and 6 of $\varepsilon_{3}=9\%$ and $\varepsilon_{6}=19\%$.
Contrarily to what is observed for the dual stripe enhancer 2+7, the correct functioning of the 3+6 enhancer relies on a higher concentration of Hb versus Kni ($\alpha_{Hb}=0.0440$ and $\alpha_{2}=0.3160$). Thus, despite depending on the same regulators as the 2+7, enhancer 3+6 has a greater affinity or a greater number of binding sites for Hb than for Kni. Note that this relationship between the ftz’s 2+7 and 3+6 enhancers is the same as between eve’s 3+7 and 4+6 enhancers. If these two dual stripe enhancers are indeed involved in Ftz’s pattern formation, it would be interesting to determine experimentally the number of binding sites in each and compare them to the calibration $\alpha$’s we computationally predicted.
We proved that stripe-specific elements can partially define the Ftz pattern along the anteroposterior axis of the Drosophila’s embryo, with stripe 4 being the only one that can’t be explained with gap transcription factors.
Figure 8: Best fits of the experimental data of the full Eve and Ftz patterns at time class 8 of mitotic cycle 14 (14A8) of Drosophila embryogenesis. These predictions were obtained following a stripe-specific enhancer theory. No regulators were identified for Ftz’s stripe 4.
3.4 Final predictions for Eve and Ftz patterns
Figure 9: Simulation of direct repression between Eve and Ftz as the main mechanism to develop the Ftz complementary pattern. Three scenarios were tested. Eve’s pattern is primarily formed by its five non-terminal enhancers, and Ftz, by some unknown process, starts developing with the (top left) same or (bottom left) similar pattern to Eve’s. Epistasis quickly overrides this mechanism and becomes the main process, creating the two complementary patterns. (bottom right) Ftz starts to be produced from its stripe-specific enhancers, and then epistasis overrides and becomes the main mechanism. We assume the ftz enhancers react to eve’s enhancers parameters. For example, we could have as an initial condition the 1+5 enhancer giving rise to a stripe overlapping Eve’s stripe 1, and the 2+7 and 3+6 enhancers originating stripes coinciding with Eve stripe pairs 3+7 and 4+6. Parameter values: (top left) k1 = 3.54, k2 = 3.54, k3 = 0.1, k4 = 0.1, k5 = 2.22, k-5 = 5.00 $\times$ 10-4, k6 = 3.66, k-6 = 5.00 $\times$ 10-4, k7 = 2.12, k-7 = 5.00 $\times$ 10-4, k8 = 2.12, k-8 = 5.00 $\times$ 10-4, $\alpha_{eve}$ = 0.365, $\alpha_{ftz}$ = 0.478; (bottom left) k1 = 3.54, k2 = 3.54, k3 = 0.1, k4 = 0.1, k5 = 2.22, k-5 = 5.00 $\times$ 10-4, k6 = 3.66, k-6 = 5.00 $\times$ 10-4, k7 = 2.12, k-7 = 5.00 $\times$ 10-4, k8 = 2.12, k-8 = 5.00 $\times$ 10-4, $\alpha_{eve}$ = 0.340, $\alpha_{ftz}$ = 0.478; (bottom right) k1 = 3.54, k2 = 3.54, k3 = 0.1, k4 = 0.1, k5 = 1.40, k-5 = 5.00 $\times$ 10-4, k6 = 3.00, k-6 = 5.00 $\times$ 10-4, k7 = 2.12, k-7 = 5.00 $\times$ 10-4, k8 = 2.12, k-8 = 5.00 $\times$ 10-4, $\alpha_{eve}$ = 0.340, $\alpha_{ftz}$ = 0.796. Suppose we reverse each protein’s contribution and the rate constants responsible for direct repression between Eve and Ftz. In that case, we obtain swapped protein concentration profiles in space (top right). In the latter case the following parameter values were considered: k1 = 3.54, k2 = 3.54, k3 = 0.1, k4 = 0.1, k5 = 3.52, k-5 = 5.00 $\times$ 10-4, k6 = 2.34, k-6 = 5.00 $\times$ 10-4, k7 = 2.12, k-7 = 5.00 $\times$ 10-4, k8 = 2.12, k-8 = 5.00 $\times$ 10-4, $\alpha_{eve}$ = 0.452, $\alpha_{ftz}$ = 0.358.
Modelling Eve and Ftz pattern development with a purely repressive gap regulating system was successful, with the two proteins demonstrating the desired complementarity. In Figure 8, we combine the two distributions obtained in Figures 4 and 7. We provided a new mechanism for the regulation of stripe 1, with a repressive combination of five gap proteins, making us reconsider the Hb activation proposal by Fujioka et al., (1999).
The final Eve and Ftz patterns have $\chi^{2}$ values of $\sqrt{\chi^{2}_{Eve}}=47$ and $\sqrt{\chi^{2}_{Ftz}}=46$, which are dependent on the system of units of experimental data. On the other hand, this depends on the variability of embryo lengths and the experimental conditions used to obtain each protein profile. These variabilities, although small, can introduce imprecisions in the stripe location predictions. Ideally, we would have to use profiles belonging to the same embryo to be objective.
3.5 An alternative Ftz pattern formation via eve-ftz epistasis
In the models developed stripe 4 expressions cannot be explained by gap transcriptional repression. The two patterns we obtained (Figure 8) are not perfectly complementary. This suggests yet another level of complexity as, for example, the introduction of cross-repression between Eve and Ftz to adjust the precise position of the stripes and develop Ftz’s stripe 4. This hypothesis advanced by Frash et al., (1987), together with studies by Lim et al., (2018) on the temporal dynamics of the two proteins, led us to consider the possibility of another, more straightforward mechanism that can override the stripe-specific enhancers action where Eve and Ftz directly repress each other’s transcription.
To test this self-repression hypothesis, we introduced the simplified kinetic mechanism
$\begin{array}[]{ll}\displaystyle\text{G}_{\text{e}}\mathop{\longrightarrow}^{k_{1}}\text{G}_{\text{e}}+\text{Eve},&\displaystyle\text{G}_{\text{f}}\mathop{\longrightarrow}^{k_{2}}\text{G}_{\text{f}}+\text{Ftz},\\ \displaystyle\text{Eve}\mathop{\longrightarrow}^{k_{3}},&\displaystyle\text{Ftz}\mathop{\longrightarrow}^{k_{4}},\\ \displaystyle\text{Eve}+\text{G}_{\text{f}}\mathop{\rightleftarrows}^{k_{5}}_{k_{-5}}\text{G}_{\text{R1}},&\displaystyle\text{Ftz}+\text{G}_{\text{e}}\mathop{\rightleftarrows}^{k_{6}}_{k_{-6}}\text{G}_{\text{R2}},\\ \displaystyle\text{TR}+\text{G}_{\text{e}}\mathop{\rightleftarrows}^{k_{7}}_{k_{-7}}\text{G}_{\text{R3}},&\displaystyle\text{TR}+\text{G}_{\text{f}}\mathop{\rightleftarrows}^{k_{8}}_{k_{-8}}\text{G}_{\text{R4}}\end{array}$ (8)
and we mimicked the stripped patterns using sums of Gaussian functions, each centred in individual stripes. The terminal repression is ensured by the hypothetical terminal regulator protein TR and was mimicked using the Error function. The evolution equations of the kinetic mechanism (8) are shown in B.
In Figure 9, we show the steady states obtained with the evolution equations associated with kinetic mechanism (8) for three different biological scenarios. In every case, this epistasis mechanism proved sufficient to produce two perfectly complementary stripe patterns. Although successful in theory, this epistasis concept implies that the Ftz pattern formation depends on Eve. At first glance, this does not agree with the mutation experiments (Harding et al., (1986)). However, we point out that there are few experiments of this type and little information on the regulation of ftz for a conclusive study. Moreover, we must not forget that eve and ftz are part of a complex gene expression network. We can conjecture a cross-repression network between the pair-rule such that other genes similar to Eve assume its role in providing periodic input in the case of a gene mutation.
3.6 Modeling segment-polarity pattern formation
Our modelling strategy can be extended to infer the behaviour of other genes in the Drosophila embryogenesis, just as the segment-polarity gene Wg. Wg has 14 stripes perfectly developed in the narrow stripes of cells that separate Eve and Ftz (Ingham et al., (1988)). We built a simple model
$\begin{array}[]{ll}\displaystyle\text{G}_{wg}\mathop{\longrightarrow}^{k_{1}}\text{G}_{wg}+\text{Wg},&\displaystyle\text{Wg}\mathop{\longrightarrow}^{k_{2}},\\ \displaystyle\text{Eve}+\text{G}_{wg}\mathop{\rightleftarrows}^{k_{3}}_{k_{-3}}\text{G}_{R1},&\displaystyle\text{Ftz}+\text{G}_{wg}\mathop{\rightleftarrows}^{k_{4}}_{k_{-4}}\text{G}_{R2}\\ \displaystyle\text{TR}+\text{G}_{wg}\mathop{\rightleftarrows}^{k_{5}}_{k_{-5}}\text{G}_{R3}.&\end{array}$ (9)
where $\text{G}_{wg}$ is the gene for the free production of the protein Wg, and Eve and Ftz are repressive regulators. As we don’t know which proteins regulate terminal gene expression, we introduced terminal repressive action by the terminal regulator TR, whose concentration distribution throughout the embryo is simulated by an Error function. The patterns of Eve and Ftz were simulated, summing localised Gaussian functions, each centred in an individual stripe. The corresponding evolution equations derived from the mass action law are shown in C. The model results in Figure 10 show the full development of the 14 stripes. By changing the parameters of the Error function, that is, its shape and the way it intersects the Eve and Ftz stripe patterns, we show how the fourteenth Wg stripe can appear before the first Eve stripe (Figure 10-(top)) or after the last Ftz stripe (Figure 10-(bottom)).
Figure 10: Simulation of a segment-polarity pattern (solid black line), through the repressive action of Eve (solid green line) and Ftz (solid red line) in Drosophila embryogenesis. We observe the clear formation of the expected 14 stripes. Terminal repression was introduced via an Error function, whose parameters determine the positioning of the fourteenth stripe: (top) before the first Eve stripe or (bottom) after the last Ftz stripe. Steady-state solutions evaluated at time $t=100$. Parameter values: (top) k1 = 13.65, k2 = 1.00, k3/4 = 5.00, k-3/-4 = 10-2, $\alpha_{0}$ = 1.00, $\alpha_{1}$ = 0.57, $\alpha_{2}$ = 0.60 and $\alpha_{4}$ = 0.15; (bottom) k1 = 13.65, k2 = 1.00, k3/4 = 5.00, k-3/-4 = 10-2, $\alpha_{0}$ = 1.00, $\alpha_{1}$ = 0.57, $\alpha_{2}$ = 0.60 and $\alpha_{4}$ = 0.42.
This demonstrates that with access to experimental data for segment-polarity expression patterns, it would be easy to analyse the possible pair-rule combinations, determine transcriptional activators and repressors for this last hierarchical level of gene regulation, and calibrate the respective patterns, regardless of the embryo in question.
4 Conclusions
We studied the gene regulatory network of Drosophila embryogenesis, focusing on modelling and calibrating the seven-stripe patterns of the proteins Eve and Ftz. We modelled seven stripe-specific enhancers for eve with the kinetic mass-action law describing transcriptional regulation. We proposed a purely repressive gap system with novel minimal regulator combinations as depicted in Figure 1. Our model predictions matched the data remarkably well. They provided a complete description of Eve’s development, including the definition of the anterior and posterior Drosophila embryo poles and the calibration of the Eve production rate across the embryo. Moreover, the calibrated model reproduces gap gene mutation experiments.
We identified three gap repressive combinations driving three ftz enhancers and successfully achieved Ftz’s pattern development calibration, except stripe 4, Figure 6. We explored an alternative mechanism of epistasis between eve and ftz that proved to be sufficient to originate two perfectly complementary patterns, having thus identified two different mechanisms for Ftz formation. Lastly, extended work inferring the Wingless (Wg) fourteen stripe pattern from Eve and Ftz repressive transcription factors have been proposed, clarifying the hierarchical structure of Drosphila’s genetic expression network during early development before cellularisation.
Ackowledgements:
We wish to express our gratitude to the two anonymous referees of the paper for their valuable comments, which significantly contributed to improving the text.
CRediT author statement
Catarina Dias: Conceptualization, Data curation, Formal analysis, Investigation, Methodology, Software, Visualization, Writing - original draft, Writing - review & editing. Rui Dilão: Conceptualization, Formal analysis, Investigation, Methodology, Software, Supervision, Writing - review & editing.
Declaration of interest
The authors declare no conflict or competing interests.
Appendix A Stripe-specific eve enhancers
A.1 Transcriptional regulation model for stripes 1 and 2 of Eve, and dual stripe 1+5 of Ftz
$\begin{array}[]{ll}\displaystyle\text{G}\mathop{\longrightarrow}^{k_{1}}\text{G}+\text{Eve},&\displaystyle\text{Eve}\mathop{\longrightarrow}^{d},\\ \displaystyle\text{Kr}+\text{G}\mathop{\rightleftarrows}^{k_{3}}_{k_{-3}}\text{G}_{R1},&\displaystyle\text{Kni}+\text{G}\mathop{\rightleftarrows}^{k_{4}}_{k_{-4}}\text{G}_{R2},\\ \displaystyle\text{Gt}+\text{G}\mathop{\rightleftarrows}^{k_{5}}_{k_{-5}}\text{G}_{R3},&\displaystyle\text{Tll}+\text{G}\mathop{\rightleftarrows}^{k_{6}}_{k_{-6}}\text{G}_{R4},\\ \displaystyle\text{Hkb}+\text{G}\mathop{\rightleftarrows}^{k_{7}}_{k_{-7}}\text{G}_{R5}.&\end{array}$ (10)
In the case of Eve we considered the gene $\text{G}\to\text{G}_{e}$ and, in the case of Ftz, $\text{G}\to\text{G}_{f}$. The stripes 1 and 2 production models of Eve and the production model for dual stripe 1+5 have different rates.
The production of proteins Eve or Ftz described by the kinetic equations (10) are
$\begin{array}[]{lcl}\overline{\text{Eve}}^{\prime}(t)&=&-d\overline{\text{Eve}}(t)+\alpha_{1}k_{1}F(\cdot)\\ \overline{\text{Kr}}^{\prime}(t)&=&k_{-3}\left[\alpha_{2}\text{G}_{R1}(0)-\overline{\text{Kr}}(t)+\overline{\text{Kr}}(0)\right]-k_{3}\overline{\text{Kr}}(t)F(\cdot)\\[5.0pt] \overline{\text{Kni}}^{\prime}(t)&=&k_{-4}\left[\alpha_{3}\text{G}_{R2}(0)-\overline{\text{Kni}}(t)+\overline{\text{Kni}}(0)\right]-k_{4}\overline{\text{Kni}}(t)F(\cdot)\\[5.0pt] \overline{\text{Gt}}^{\prime}(t)&=&k_{-5}\left[\alpha_{4}\text{G}_{R3}(0)-\overline{\text{Gt}}(t)+\overline{\text{Gt}}(0)\right]-k_{5}\overline{\text{Gt}}(t)F(\cdot)\\[5.0pt] \overline{\text{Tll}}^{\prime}(t)&=&k_{-6}\left[\alpha_{6}\text{G}_{R4}(0)-\overline{\text{Tll}}(t)+\overline{\text{Tll}}(0)\right]-k_{6}\overline{\text{Tll}}(t)F(\cdot)\\[5.0pt] \overline{\text{Hkb}}^{\prime}(t)&=&k_{-7}\left[\alpha_{7}\text{G}_{R5}(0)-\overline{\text{Hkb}}(t)+\overline{\text{Hkb}}(0)\right]-k_{7}\overline{\text{Hkb}}(t)F(\cdot),\end{array}$
where
$\begin{array}[]{ll}F(\cdot)=&\left(\frac{\overline{\text{G}}(0)}{\alpha_{0}}+\frac{\overline{\text{Kr}}(t)-\overline{\text{Kr}}(0)}{\alpha_{2}}+\frac{\overline{\text{Kni}}(t)-\overline{\text{Kni}}(0)}{\alpha_{3}}+\frac{\overline{\text{Gt}}(t)-\overline{\text{Gt}}(0)}{\alpha_{4}}\right.\\ &\left.+\frac{\overline{\text{Tll}}(t)-\overline{\text{Tll}}(0)}{\alpha_{6}}+\frac{\overline{\text{Hkb}}(t)-\overline{\text{Hkb}}(0)}{\alpha_{7}}\right),\end{array}$
P or $=\alpha_{\text{p}}\text{P}$ P represent a specific protein. The calibration parameters for each repressor proteins are: $\alpha_{0}=\alpha_{G_{e}/G_{f}}$, $\alpha_{1}=\alpha_{eve/ftz}$, $\alpha_{2}=\alpha_{kr}$, $\alpha_{3}=\alpha_{kni}$, $\alpha_{4}=\alpha_{gt}$, $\alpha_{5}=\alpha_{hb}$, $\alpha_{6}=\alpha_{tll}$ and $\alpha_{7}=\alpha_{hkb}$. In all the simulations, we have assumed $\alpha_{0}=1$. These conditions are fixed for all the models analysed.
At each anteroposterior $x$-coordinate of the embryo, the repressive protein data are introduced at $\overline{\text{P}}(0)$. The initial conditions for each repressor complex are $\text{G}_{R1}(0)=\ldots=\text{G}_{R5}(0)=0$ and $\overline{\text{G}}(0)=250$. The initial conditions for Eve and Ftz models are $\overline{\text{Eve}}(0)=0$ and $\overline{\text{Ftz}}(0)=0$.
A.2 Transcriptional regulation model for dual stripes 3+7 of Eve, and dual stripe 2+7 and 3+6 of Ftz
$\begin{array}[]{ll}\displaystyle\text{G}\mathop{\longrightarrow}^{k_{1}}\text{G}+\text{Eve},&\displaystyle\text{Eve}\mathop{\longrightarrow}^{d},\\ \displaystyle\text{Kni}+\text{G}\mathop{\rightleftarrows}^{k_{3}}_{k_{-3}}\text{G}_{R1},&\displaystyle\text{Hb}+\text{G}\mathop{\rightleftarrows}^{k_{4}}_{k_{-4}}\text{G}_{R2},\\ \displaystyle\text{Tll}+\text{G}\mathop{\rightleftarrows}^{k_{5}}_{k_{-5}}\text{G}_{R3},&\displaystyle\text{Hkb}+\text{G}\mathop{\rightleftarrows}^{k_{6}}_{k_{-6}}\text{G}_{R4}.\end{array}$ (11)
The production of protein Eve or Ftz described by the kinetic equations (11) are
$\begin{array}[]{lcl}\overline{\text{Eve}}^{\prime}(t)&=&-d\overline{\text{Eve}}(t)+\alpha_{1}k_{1}F(\cdot)\\[5.0pt] \overline{\text{Kni}}^{\prime}(t)&=&k_{-3}\left[\alpha_{3}\text{G}_{R1}(0)-\overline{\text{Kni}}(t)+\overline{\text{Kni}}(0)\right]-k_{3}\overline{\text{Kni}}(t)F(\cdot)\\[5.0pt] \overline{\text{Hb}}^{\prime}(t)&=&k_{-4}\left[\alpha_{5}\text{G}_{R2}(0)-\overline{\text{Hb}}(t)+\overline{\text{Hb}}(0)\right]-k_{4}\overline{\text{Hb}}(t)F(\cdot)\\[5.0pt] \overline{\text{Tll}}^{\prime}(t)&=&k_{-5}\left[\alpha_{6}\text{G}_{R3}(0)-\overline{\text{Tll}}(t)+\overline{\text{Tll}}(0)\right]-k_{5}\overline{\text{Tll}}(t)F(\cdot)\\[5.0pt] \overline{\text{Hkb}}^{\prime}(t)&=&k_{-6}\left[\alpha_{7}\text{G}_{R4}(0)-\overline{\text{Hkb}}(t)+\overline{\text{Hkb}}(0)\right]-k_{6}\overline{\text{Hkb}}(t)F(\cdot),\end{array}$
where
$\begin{array}[]{ll}F(\cdot)=&\left(\frac{\overline{\text{G}}(0)}{\alpha_{0}}+\frac{\overline{\text{Kni}}(t)-\overline{\text{Kni}}(0)}{\alpha_{3}}+\frac{\overline{\text{Hb}}(t)-\overline{\text{Hb}}(0)}{\alpha_{5}}+\frac{\overline{\text{Tll}}(t)-\overline{\text{Tll}}(0)}{\alpha_{6}}\right.\\ &\left.+\frac{\overline{\text{Hkb}}(t)-\overline{\text{Hkb}}(0)}{\alpha_{7}}\right).\end{array}$
A.3 Transcriptional regulation model for dual stripes 4+6 of Eve
$\begin{array}[]{ll}\displaystyle\text{G}_{e}\mathop{\longrightarrow}^{k_{1}}\text{G}_{e}+\text{Eve},&\displaystyle\text{Eve}\mathop{\longrightarrow}^{d},\\ \displaystyle\text{Kni}+\text{G}_{e}\mathop{\rightleftarrows}^{k_{3}}_{k_{-3}}\text{G}_{R1},&\displaystyle\text{Gt}+\text{G}_{e}\mathop{\rightleftarrows}^{k_{4}}_{k_{-4}}\text{G}_{R2},\\ \displaystyle\text{Hb}+\text{G}_{e}\mathop{\rightleftarrows}^{k_{5}}_{k_{-5}}\text{G}_{R3},&\displaystyle\text{Tll}+\text{G}_{e}\mathop{\rightleftarrows}^{k_{6}}_{k_{-6}}\text{G}_{R4},\\ \displaystyle\text{Hkb}+\text{G}_{e}\mathop{\rightleftarrows}^{k_{7}}_{k_{-7}}\text{G}_{R5}.&\end{array}$ (12)
The production of protein Eve described by the kinetic equations (12) is
$\begin{array}[]{lcl}\overline{\text{Eve}}^{\prime}(t)&=&-d\overline{\text{Eve}}(t)+\alpha_{1}k_{1}F(\cdot)\\ \overline{\text{Kni}}^{\prime}(t)&=&k_{-3}\left[\alpha_{3}\text{G}_{R1}(0)-\overline{\text{Kni}}(t)+\overline{\text{Kni}}(0)\right]-k_{3}\overline{\text{Kni}}(t)F(\cdot)\\[5.0pt] \overline{\text{Gt}}^{\prime}(t)&=&k_{-4}\left[\alpha_{4}\text{G}_{R2}(0)-\overline{\text{Gt}}(t)+\overline{\text{Gt}}(0)\right]-k_{4}\overline{\text{Gt}}(t)F(\cdot)\\[5.0pt] \overline{\text{Hb}}^{\prime}(t)&=&k_{-5}\left[\alpha_{5}\text{G}_{R3}(0)-\overline{\text{Hb}}(t)+\overline{\text{Hb}}(0)\right]-k_{5}\overline{\text{Hb}}(t)F(\cdot)\\[5.0pt] \overline{\text{Tll}}^{\prime}(t)&=&k_{-6}\left[\alpha_{6}\text{G}_{R4}(0)-\overline{\text{Tll}}(t)+\overline{\text{Tll}}(0)\right]-k_{6}\overline{\text{Tll}}(t)F(\cdot)\\[5.0pt] \overline{\text{Hkb}}^{\prime}(t)&=&k_{-7}\left[\alpha_{7}\text{G}_{R5}(0)-\overline{\text{Hkb}}(t)+\overline{\text{Hkb}}(0)\right]-k_{7}\overline{\text{Hkb}}(t)F(\cdot),\end{array}$
where
$\begin{array}[]{ll}F(\cdot)=&\left(\frac{\overline{\text{G}}_{e}(0)}{\alpha_{0}}+\frac{\overline{\text{Kni}}(t)-\overline{\text{Kni}}(0)}{\alpha_{3}}+\frac{\overline{\text{Gt}}(t)-\overline{\text{Gt}}(0)}{\alpha_{4}}+\frac{\overline{\text{Hb}}(t)-\overline{\text{Hb}}(0)}{\alpha_{5}}\right.\\ &\left.+\frac{\overline{\text{Tll}}(t)-\overline{\text{Tll}}(0)}{\alpha_{6}}+\frac{\overline{\text{Hkb}}(t)-\overline{\text{Hkb}}(0)}{\alpha_{7}}\right).\end{array}$
A.4 Transcriptional regulation model for stripe 5 of Eve
$\begin{array}[]{ll}\displaystyle\text{G}_{e}\mathop{\longrightarrow}^{k_{1}}\text{G}_{e}+\text{Eve},&\displaystyle\text{Eve}\mathop{\longrightarrow}^{d},\\ \displaystyle\text{Kr}+\text{G}_{e}\mathop{\rightleftarrows}^{k_{3}}_{k_{-3}}\text{G}_{R1},&\displaystyle\text{Gt}+\text{G}_{e}\mathop{\rightleftarrows}^{k_{4}}_{k_{-4}}\text{G}_{R2},\\ \displaystyle\text{Hb}+\text{G}_{e}\mathop{\rightleftarrows}^{k_{5}}_{k_{-5}}\text{G}_{R3},&\displaystyle\text{Tll}+\text{G}_{e}\mathop{\rightleftarrows}^{k_{6}}_{k_{-6}}\text{G}_{R4},\\ \displaystyle\text{Hkb}+\text{G}_{e}\mathop{\rightleftarrows}^{k_{7}}_{k_{-7}}\text{G}_{R5}.&\end{array}$ (13)
The production of protein Eve described by the kinetic equations (13) is
$\begin{array}[]{lcl}\overline{\text{Eve}}^{\prime}(t)&=&-d\overline{\text{Eve}}(t)+\alpha_{1}k_{1}F(\cdot)\\ \overline{\text{Kr}}^{\prime}(t)&=&k_{-3}\left[\alpha_{2}\text{G}_{R1}(0)-\overline{\text{Kr}}(t)+\overline{\text{Kr}}(0)\right]-k_{3}\overline{\text{Kr}}(t)F(\cdot)\\[5.0pt] \overline{\text{Gt}}^{\prime}(t)&=&k_{-4}\left[\alpha_{4}\text{G}_{R2}(0)-\overline{\text{Gt}}(t)+\overline{\text{Gt}}(0)\right]-k_{4}\overline{\text{Gt}}(t)F(\cdot)\\[5.0pt] \overline{\text{Hb}}^{\prime}(t)&=&k_{-5}\left[\alpha_{5}\text{G}_{R3}(0)-\overline{\text{Hb}}(t)+\overline{\text{Hb}}(0)\right]-k_{5}\overline{\text{Hb}}(t)F(\cdot)\\[5.0pt] \overline{\text{Tll}}^{\prime}(t)&=&k_{-6}\left[\alpha_{6}\text{G}_{R4}(0)-\overline{\text{Tll}}(t)+\overline{\text{Tll}}(0)\right]-k_{6}\overline{\text{Tll}}(t)F(\cdot)\\[5.0pt] \overline{\text{Hkb}}^{\prime}(t)&=&k_{-7}\left[\alpha_{7}\text{G}_{R5}(0)-\overline{\text{Hkb}}(t)+\overline{\text{Hkb}}(0)\right]-k_{7}\overline{\text{Hkb}}(t)F(\cdot),\end{array}$
where
$\begin{array}[]{ll}F(\cdot)=&\left(\frac{\overline{\text{G}}_{e}(0)}{\alpha_{0}}+\frac{\overline{\text{Kr}}(t)-\overline{\text{Kr}}(0)}{\alpha_{2}}+\frac{\overline{\text{Gt}}(t)-\overline{\text{Gt}}(0)}{\alpha_{4}}+\frac{\overline{\text{Hb}}(t)-\overline{\text{Hb}}(0)}{\alpha_{5}}\right.\\ &\left.+\frac{\overline{\text{Tll}}(t)-\overline{\text{Tll}}(0)}{\alpha_{6}}+\frac{\overline{\text{Hkb}}(t)-\overline{\text{Hkb}}(0)}{\alpha_{7}}\right).\end{array}$
Appendix B Alternative pattern formation: eve-ftz epistasis
$\begin{array}[]{ll}\displaystyle\text{G}_{\text{e}}\mathop{\longrightarrow}^{k_{1}}\text{G}_{\text{e}}+\text{Eve},&\displaystyle\text{G}_{\text{f}}\mathop{\longrightarrow}^{k_{2}}\text{G}_{\text{f}}+\text{Ftz},\\ \displaystyle\text{Eve}\mathop{\longrightarrow}^{k_{3}},&\displaystyle\text{Ftz}\mathop{\longrightarrow}^{k_{4}},\\ \displaystyle\text{Eve}+\text{G}_{\text{f}}\mathop{\rightleftarrows}^{k_{5}}_{k_{-5}}\text{G}_{\text{R1}},&\displaystyle\text{Ftz}+\text{G}_{\text{e}}\mathop{\rightleftarrows}^{k_{6}}_{k_{-6}}\text{G}_{\text{R2}},\\ \displaystyle\text{TR}+\text{G}_{\text{e}}\mathop{\rightleftarrows}^{k_{7}}_{k_{-7}}\text{G}_{\text{R3}},&\displaystyle\text{TR}+\text{G}_{\text{f}}\mathop{\rightleftarrows}^{k_{8}}_{k_{-8}}\text{G}_{\text{R4}}.\end{array}$ (14)
The production of proteins Eve and Ftz described by the kinetic equations (14) is
$\begin{array}[]{lcl}\overline{\text{Eve}}^{\prime}(t)&=&-\overline{\text{Eve}}(t)\left(k_{3}+\frac{\overline{\text{G}}_{f}(t)k_{5}}{\alpha_{1}}\right)+\frac{\overline{\text{G}}_{e}(t)k_{1}\alpha_{2}}{\alpha_{0}}+k_{-5}\alpha_{2}F_{1}(\cdot)\\[5.0pt] \overline{\text{Ftz}}^{\prime}(t)&=&-\overline{\text{Ftz}}(t)\left(k_{4}+\frac{\overline{\text{G}}_{e}(t)k_{6}}{\alpha_{0}}\right)+\alpha_{3}\left(k_{-6}\text{G}_{R2}(t)+\frac{\overline{G}_{f}(t)k_{2}}{\alpha_{1}}\right)\\[5.0pt] \overline{\text{G}}_{e}^{\prime}(t)&=&-\overline{\text{G}}_{e}(t)\left(\frac{\overline{\text{Ftz}}(t)k_{6}}{\alpha_{3}}+\frac{\overline{\text{TR}}(t)k_{7}}{\alpha_{4}}\right)+k_{-6}\alpha_{0}\text{G}_{R2}(t)-k_{-7}\alpha_{0}F_{2}(\cdot)\\[5.0pt] \overline{\text{G}}_{f}^{\prime}(t)&=&-\overline{\text{G}}_{f}(t)\left(\frac{\overline{\text{Eve}}(t)k_{5}}{\alpha_{2}}+\frac{\overline{\text{TR}}(t)k_{8}}{\alpha_{4}}\right)+k_{-5}\alpha_{1}F_{1}(\cdot)+k_{-8}\alpha_{1}F_{3}(\cdot)\\[5.0pt] \overline{\text{TR}}^{\prime}(t)&=&-\overline{\text{TR}}(t)\left(\frac{\overline{\text{G}}_{e}(t)k_{7}}{\alpha_{0}}+\frac{\overline{\text{G}}_{f}(t)k_{8}}{\alpha_{1}}\right)-k_{-7}\alpha_{4}F_{2}(\cdot)+k_{-8}\alpha_{4}F_{3}(\cdot)\\[5.0pt] \text{G}_{R2}^{\prime}(t)&=&-\text{G}_{R2}(t)k_{-6}+\frac{\overline{\text{Ftz}}(t)\overline{G}_{e}(t)k_{6}}{\alpha_{0}\alpha_{3}},\end{array}$
where
$\begin{array}[]{lcl}F_{1}(\cdot)&=&\left(\text{G}_{R1}(0)+\text{G}_{R2}(0)-\text{G}_{R2}(t)+\frac{\overline{\text{G}}_{e}(0)-\overline{\text{G}}_{e}(t)}{\alpha_{0}}+\frac{\overline{\text{G}}_{f}(0)-\overline{\text{G}}_{f}(t)}{\alpha_{1}}\right.\\ &&+\left.\frac{\overline{\text{TR}}(t)-\overline{\text{TR}}(0)}{\alpha_{4}}\right)\\[5.0pt] F_{2}(\cdot)&=&\left(\text{G}_{R2}(0)-\text{G}_{R3}(0)-\text{G}_{R2}(t)+\frac{\overline{\text{G}}_{f}(0)-\overline{\text{G}}_{f}(t)}{\alpha_{1}}+\frac{\overline{\text{TR}}(t)-\overline{\text{TR}}(0)}{\alpha_{4}}\right)\\[5.0pt] F_{3}(\cdot)&=&\left(\text{G}_{R2}(0)+\text{G}_{R4}(0)-\text{G}_{R2}(t)+\frac{\overline{\text{G}}_{f}(0)-\overline{\text{G}}_{f}(t)}{\alpha_{1}}\right).\end{array}$
The calibration parameter parameters are: $\alpha_{0}=\alpha_{G_{e}}=1$, $\alpha_{1}=\alpha_{G_{f}}=1$, $\alpha_{2}=\alpha_{eve}$, $\alpha_{3}=\alpha_{ftz}$, $\alpha_{4}=\alpha_{tr}$.
Appendix C Segment-Polarity pattern formation
$\begin{array}[]{ll}\displaystyle\text{G}_{wg}\mathop{\longrightarrow}^{k_{1}}\text{G}_{wg}+\text{Wg},&\displaystyle\text{Wg}\mathop{\longrightarrow}^{k_{2}},\\ \displaystyle\text{Eve}+\text{G}_{wg}\mathop{\rightleftarrows}^{k_{3}}_{k_{-3}}\text{G}_{R1},&\displaystyle\text{Ftz}+\text{G}_{wg}\mathop{\rightleftarrows}^{k_{4}}_{k_{-4}}\text{G}_{R2}\\ \displaystyle\text{TR}+\text{G}_{wg}\mathop{\rightleftarrows}^{k_{5}}_{k_{-5}}\text{G}_{R3}.&\end{array}$ (15)
The production of proteins Eve, Ftz and Wg described by the kinetic equations (15) is
$\begin{array}[]{lcl}\overline{\text{Eve}}^{\prime}(t)&=&k_{-3}\left[\alpha_{1}\text{G}_{R1}(0)-\overline{\text{Eve}}(t)+\overline{\text{Eve}}(0)\right]+k_{3}\overline{\text{Eve}}(t)F(\cdot)\\[5.0pt] \displaystyle\overline{\text{Ftz}}^{\prime}(t)&=&k_{-4}\left[\alpha_{2}\text{G}_{R2}(0)-\overline{\text{Ftz}}(t)+\overline{\text{Ftz}}(0)\right]+k_{4}\overline{\text{Ftz}}(t)F(\cdot)\\[5.0pt] \displaystyle\overline{\text{TR}}^{\prime}(t)&=&k_{-5}\left[\alpha_{3}\text{G}_{R3}(0)-\overline{\text{TR}}(t)+\overline{\text{TR}}(0)\right]-k_{5}\overline{\text{TR}}(t)F(\cdot)\\[5.0pt] \displaystyle\overline{\text{Wg}}^{\prime}(t)&=&\alpha_{4}k_{1}F(\cdot)-k_{2}\overline{\text{Wg}}(t),\end{array}$
where
$F(\cdot)=\left(\frac{\overline{\text{Eve}}(0)-\overline{\text{Eve}}(t)}{\alpha_{1}}+\frac{\overline{\text{Ftz}}(0)-\overline{\text{Ftz}}(t)}{\alpha_{2}}+\frac{\overline{\text{TR}}(0)-\overline{\text{TR}}(t)}{\alpha_{3}}-\frac{\overline{\text{G}}_{wg}(0)}{\alpha_{0}}\right).$
The calibration parameter parameter are: $\alpha_{0}=\alpha_{G_{wg}}=1$, $\alpha_{1}=\alpha_{eve}$, $\alpha_{2}=\alpha_{ftz}$, $\alpha_{3}=\alpha_{tr}$, $\alpha_{4}=\alpha_{wg}=1$.